Calculate the number of pounds of CO2 released into the atmosphere when a 14.0 gallon tank of gasoline is burned in an automobile engine. Assume that gasoline is primarily octane, C8H18, and that the density of gasoline is 0.692 g⋅mL−1. This assumption ignores additives. Also, assume complete combustion.
1
Expert's answer
2020-07-07T14:36:08-0400

14.0 gallon tank contains 14.0·3.785=52.99 L of gasoline. Making use of the density, we calculate the mass of the gasoline:

"m = d·V = 0.692(\text{g/mL)}·52.99·10^3(\text{mL}) = 36.67·10^3" g.

As the gasoline is primarily octane, the mass of octane is 36.67·103 g. The number of the moles of octane is its mass divided by its molar mass "M = 114.23" g/mol:

"n = \frac{m}{M} = \frac{36.67·10^3 (g)}{114.23 (g/mol)} = 321" mol.

The equation of the reaction between octane and oxygen is:

2C8H18 + 25O2 "\rightarrow" 16CO2 + 18H2O.

As one can see, when 2 mol of octane reacts, 16 mol of CO2 is produced:

"\frac{n(C_8H_{18})}{2} = \frac{n(CO_2)}{16}".

Therefore, the number of the moles of CO2 produced is:

"n(CO_2) = 321\u00b716\/2 = 2568" mol.

The molar mass of CO2 is 44.01 g/mol. Thus:

"m(CO_2) = n\u00b7M =2568\u00b744.01 = 1.13\u00b710^{5}" g.

In pounds, the mass of CO2 will be:

"\frac{1.13·10^5}{453.59} = 249" lb.

Answer: 249 lb of CO2 is released into the atmosphere when a 14.0 gallon tank of gasoline is burned in an automobile engine.